A draw is going to be played in 14 people. Selectors would select 6 people. Then what is the winning percentage of individual person ?
1
Expert's answer
2013-01-30T09:01:32-0500
The probability of individual person to win is

Pro = 6/14 = 3/7.

So, the winning percentage of individual person is

Per = (3/7)·100% ≈ 0.4285714·100% = 42.85714%.